# A Transformer Model for Predicting Chemical Reaction Products from Generic Templates

### v2

@@ Abstract
The accurate prediction of chemical reaction outcomes is a major challenge in computational chemistry. Current models rely heavily on either highly specific reaction templates or template-free methods, both of which present limitations. To address these limitations, this work proposes the Broad Reaction Set (BRS), a dataset featuring 20 generic reaction templates that allow for the efficient exploration of the chemical space. Additionally, ProPreT5 is introduced, a T5 model tailored to chemistry that achieves a balance between rigid templates and template-free methods. ProPreT5 demonstrates its capability to generate accurate, valid, and realistic reaction products, making it a promising solution that goes beyond the current state-of-the-art on the complex reaction product prediction task.
@@ I Introduction
The accurate prediction of chemical reaction outcomes is an important task in chemistry as it allows the construction of organic synthesis routes. Given a set of reactive molecules i.e. reactants, the goal is to determine the outcome, the product. This task is especially challenging, requiring a thorough understanding of chemical substances, compound classes, reactions, underlying reactivity patterns, and reaction conditions. Such an understanding is fundamental for various applications, including: Drug discovery [1], material science [2], and green chemistry [3].
Organic chemistry synthesis route planning has usually been an expert driven and rule based approach. However, the advancement of machine learning has transformed the field of cheminformatics. Among the most promising techniques in this field are Transformer models [4], originally developed for sequence-to-sequence tasks in Natural Language Processing (NLP) [5]. The Transformer architecture has revolutionized NLP by enabling models to understand context through self-attention mechanisms. This capability helps the models to learn to dynamically assign weights to different parts of the input data to produce more accurate and contextually relevant outputs. The Transformer architecture has also been successfully applied to the field of chemistry in various tasks, such as single-step chemical reaction prediction [6, 7, 8], retrosynthesis [8, 9, 10], molecule generation [11, 12, 13], and molecular property prediction[14, 15, 16]. However, these models are highly dependent on the quality and diversity of the data used for training.
Molecular data can be represented in the form of strings which are derived from a graph traversal of their molecular structure. This compact representation, known as the Simplified Molecular Input Line Entry System (SMILES) [17] encodes molecules where each character represents an atom or bond, providing a linear representation useful for computational purposes and database storage. When the graph traversal follows the specific order established by the notation, the resulting SMILES is canonical and a molecule has only one canonical SMILES but can have multiple non-canonical SMILES representations.
@@
Publicly available reaction datasets are mostly obtained through data extraction from patents published by The United States Patent and Trademark Office (USPTO) [19]. These reactions are represented using SMILES representation along with their reactants and reagents. Previous work has created a sub-dataset called USPTO MIT [20] by filtering it to 470,000 examples and splitting it into train/test/validation sets. USPTO MIT datasets are widely used in the literature.
Since USPTO datasets are derived from patents, they provide highly specific reactions tailored exclusively to particular reactant-product combinations. Furthermore, the chemical space explored using these datasets is limited to the information contained within the patented reactions. Consequently, we assume that relying solely on patents introduces bias during training, leaving a substantial portion of the chemical space unaccounted for. While previous work [8] has demonstrated great accuracy in predicting reaction outcomes using this dataset, the lack of molecule discovery and real-world applications stemming from such models highlights a key limitation: models trained solely on the patented chemical space suffer from generalization issues and are unsuitable for practical applications [21]. Additionally, the limitations of the existing datasets become particularly evident when reactions are considered as actions in reinforcement learning. The specificity of the reactions limits the exploration of all possible combinatorial space with these algorithms.
@@
Reaction template datasets are often argued to be difficult to maintain due to the endless and unmanageable number of reactant-product combinations [7], leading many to favor template-free methods as the future of the field. While maintaining templates for highly specific reactions, such as those in USPTO, is undeniably challenging, we argue that employing generic reactions, like those in the proposed BRS, combines the benefits of reaction templates while addressing the maintenance issues.
To investigate the properties of BRS, we release a tailored T5 [22] model, ProPreT5, designed for the reaction prediction problem. ProPreT5 was separately trained in both template-based and template-free settings using the BRS and USPTO MIT datasets. This approach led to several key insights. When using reaction templates to predict products with the USPTO MIT dataset, the task proved to be trivial, as the product patterns in the reaction templates essentially enumerate the product molecule, significantly simplifying the prediction. The goal of this study is not to improve the USPTO MIT baseline but to highlight its limitations and propose a more realistic alternative. To this end, we opted for an economical training setup, using the USPTO MIT dataset as a sanity check for ProPreT5. Ultimately, we demonstrate that ProPreT5 generates realistic, valid, and accurate reaction products while allowing for the exploration of the entire chemical space in a template-based setting, addressing key challenges and going beyond the current state-of-the-art approaches in reaction prediction. An analysis aimed at improving the interpretability of the template-based ProPreT5 model is also proposed, leading to the conclusion that generic reaction templates provide the model with crucial information for product generation.
The contribution presented in this article is threefold:
  • We address the limitations of publicly available reaction datasets by introducing a novel reaction set BRS allowing for a broader exploration of the chemical space.
  • We release ProPreT5, an improved and highly flexible T5-based model capable of generating realistic, valid, and accurate products.
  • We perform an interpretability analysis of the model to identify key contextual information that affects the accuracy of product prediction.
TABLE I: Generic Reactions Patterns
The greater-than signs (>>much-greater-than$>$ >) separates reactants from products, while a dot (.) distinguishes individual molecules. #$\#n$ represents any atom with the atomic number $n$, and specific letters indicate particular chemical environments; $:n$ is used to map and track specific subgraphs within the reaction. For more information on the SMARTS notation refer to [18].
@@
5 [#6,#7,#8;h:1$\sim$[*:2$\sim$[#6,#7,#8;h:3]>> [#6,#7,#8:1]1[*:2$\sim$[#6,#7,#8:3]1 15 [#6,#7,#8:1]1[*:2$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[#6,#7,#8;h:3]
6 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[#6,#7,#8;h:3]>> [#6,#7,#8:1]1[*:2$\sim$[*:4$\sim$[#6,#7,#8:3]1 16 [#6,#7,#8:1]1[*:2$\sim$[*:4$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[#6,#7,#8;h:3]
7 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8;h:3]>> [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8:3]1 17 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8:3]1>> [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[#6,#7,#8;h:3]
8 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$ [#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6] $\sim$ [#6,#7,#8:3]1 18 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[#6,#7,#8:3]1 >>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$ [#6,#7,#8;h:3]
9 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6] $\sim$[*:7$\sim$[#6,#7,#8:3]1 19 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [#6,#7,#8:3]1>>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[#6,#7,#8;h:3]
10 [#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$ [*:8$\sim$[#6,#7,#8;h:3]>>[O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[*:8$\sim$[#6,#7,#8:3]1 20 [O,N,C:1]1[*:2$\sim$[*:4$\sim$[*:5$\sim$[*:6$\sim$[*:7$\sim$[*:8] $\sim$[#6,#7,#8:3]1>>[#6,#7,#8;h:1$\sim$[*:2$\sim$[*:4$\sim$[*:5] $\sim$[*:6$\sim$[*:7$\sim$[*:8$\sim$[#6,#7,#8;h:3]
@@ II Related Work
Most of the recent literature on template-based approaches in reaction prediction has focused on graph-based models. [23, 24]. However, to the best of our knowledge, the potential of Transformer models for template-based sequence prediction has yet to be fully explored. This can be partly attributed to the specificity of reaction templates in the benchmark datasets (USPTO MIT).
Template-free, sequence-based models have achieved impressive results on the USPTO MIT benchmark, holding the current state-of-the-art in the single-step product prediction task [8]. However, these models face significant challenges in terms of applicability. The USPTO MIT dataset, limited to reactions found in patents, struggle to capture the diversity and complexity of real-world organic synthesis. As a result, while template-free models perform well in generating reaction outcomes, they struggle to generalize to the discovery of novel molecules or the prediction of reaction outcomes that fall outside the dataset [21].
Notable advancements in this domain include Molecular Transformer [6], which used a smaller version of the base transformers architecture [4]. The model was trained on both forward and backward tasks. This was the first application of transformers to the single-step reaction prediction task. Augmented Molformer [7] used data augmentation by extending the inputs with non-canonical SMILES and demonstrated that doing so increased the model’s generalization. Chemformer [8], on the other hand, took things even further by proposing a much larger BART [25] model with special pretraining. The pretraining included masking and data augmentation, making the model more robust and increasing prediction accuracy. While these template-free transformer models have demonstrated remarkable success in reaction prediction, the exploration of template-based sequence-to-sequence models for this task remains an underexplored area, leaving room for further development in this domain.
The lack of real-world applications using previous models highlights the need for a dataset with generic reactions, offering a more versatile training ground for reaction prediction models. This would allow models to explore the entire chemical space, ultimately advancing the field beyond the limitations of current benchmarks.
@@ III Datasets
Two datasets are employed in this work: USPTO MIT [20] and the dataset constructed by using the proposed BRS which is one of the key contributions of this work. The USPTO MIT dataset contains reactions, along with the corresponding reactants, reagents, and products extracted from patents, represented in SMILES notation. The USPTO MIT dataset includes highly specific reaction templates, valid only for the exact combination of reactants, reagents, and products. It is divided into two subsets: USPTO MIT Mixed, which does not distinguish between reactants and reagents and is considered a slightly more challenging problem, and USPTO MIT Separated, which separates reagents from reactants.
In this study, we worked with the USPTO MIT Mixed dataset and performed a verification process using RDKit [26] to ensure consistency between reactants, reagents, products, and their corresponding reaction templates. With our specific configuration, we identified a small number of discrepancies where the combination of reactants, reagents, and products did not align with the reaction template. These mismatches indicated that the reactions could not be performed as defined, leading to the removal of 142 examples from the dataset. This adjustment had a negligible impact on the overall dataset size: the training set was reduced from 409,035 to 408,916 examples, the validation set from 30,000 to 29,990, and the test set from 40,000 to 39,987. The proposed train/validation/test split was preserved.
@@ III-A Proposed Reaction Set: Broad Reaction Set
For the proposed dataset, 20 new generic reactions are introduced, as shown in Table I. Ten of these reactions are constructive, and the other ten are destructive, represented in SMARTS notation. These reactions are inspired by those used in an evolutionary algorithm named EvoMol, which demonstrated efficiency in exploring chemical space [27]. These reactions serve as foundational building blocks for constructing prediction datasets similar to USPTO. Since these reactions can be applied to a wide range of molecules, they offer significant flexibility for use with publicly available or commercial datasets.
Destructive reactions, the reverse of constructive reactions, enable a return to an earlier stage in the chemical space, offering the possibility to explore alternative pathways. Constructive and destructive reactions are symmetrical, and since this work focuses on single-step prediction, only constructive reactions are used in this study.
@@
For simplicity, the transformations defined in these reactions are currently limited to atoms C, N, O, and F. However, this does not imply that the reactions cannot be applied to molecules containing other atoms. It simply means that only the reactivity of these four atoms is considered, and transformations will occur exclusively between them within the molecule. Additionally, reactions constructing cycles up to size 8 were defined. This limitation is partly due to challenges in defining reactions that can generate cycles of arbitrary size using SMARTS notation, but also because cycles larger than size 8 are rare. Moreover, chemists have proposed metrics that try to assess the synthesizability of a molecular graph such as SAScore [28] where large cycles are penalized.
@@
III-B Construction of The Proposed Dataset
To construct the dataset used in this study, reactants were randomly sampled from several publicly available datasets: EVO10 [29], which enumerates every possible molecule with a maximum of 10 atoms of either C, N, O, F, or S; ZINC20 [30]; and ChEMBL34 [31], both of which were filtered to the same atoms as EVO10. For simplicity, the proposed dataset will be referred to as the BRS dataset, even though BRS is the name given to the reaction set employed for the construction of the dataset.
To create the dataset, molecules were randomly selected from the three datasets mentioned above while making sure that ZINC20, a dataset containing over 1B molecules, wasn’t overly represented. At each iteration, either a new molecule was selected from the datasets or a product from an earlier iteration was used as a reactant, allowing for the construction of multi-step synthesis routes for future studies. To avoid the generation of unrealistic products by the generic BRS, several filtering steps were applied. First, only canonical SMILES were kept. Additionally, filters described in [29] were implemented: one removed molecules with ECFP4 subgraphs absent from real-molecule datasets such as ChEMBL and ZINC, while the other excluded products containing Generic Cyclic Features (GCF)—scaffolds of molecular cycles not observed in the same datasets. Products that passed these filters were deemed realistic. To enhance diversity, multiple products were included for the same reactant-reaction combinations, allowing the model to learn that a single input could yield multiple valid outputs. We also ensured that each reaction appeared an equal number of times in the train, test, and validation datasets. The resulting train dataset contains 311,621 examples, while the validation and test datasets both contain 10,000 examples. Much larger datasets can be constructed by using the reactions from the BRS and by incorporating molecules from different datasets.

### v3

# A Transformer Model for Predicting Chemical Products from Generic SMARTS Templates with Data Augmentation

@@ Abstract
The accurate prediction of chemical reaction outcomes is a major challenge in computational chemistry. Current models rely heavily on either highly specific reaction templates or template-free methods, both of which present limitations. To address these, this work proposes the Broad Reaction Set (BRS), a set featuring 20 generic reaction templates written in SMARTS, a pattern-based notation designed to describe substructures and reactivity. Additionally, we introduce ProPreT5, a T5-based model specifically adapted for chemistry and, to the best of our knowledge, the first language model capable of directly handling and applying SMARTS reaction templates. To further improve generalization, we propose the first augmentation strategy for SMARTS, which injects structural diversity at the pattern level. Trained on augmented templates, ProPreT5 demonstrates strong predictive performance and generalization to unseen reactions. Together, these contributions provide a novel and practical alternative to current methods, advancing the field of template-based reaction prediction.
@@ I Introduction
The accurate prediction of chemical reaction outcomes is an important task in chemistry as it allows the construction of organic synthesis routes. Given a set of reactive molecules i.e., reactants, the goal is to determine the outcome, the product. This task is especially challenging, requiring a thorough understanding of chemical substances, compound classes, reactions, underlying reactivity patterns, and reaction conditions. Such an understanding is fundamental for various applications, including: Drug discovery [1], material science [2], and green chemistry [3].
Organic chemistry synthesis route planning has usually been an expert-driven and rule-based approach. However, the advancement of machine learning has transformed the field of cheminformatics. Among the most promising techniques in this field are Transformer models [4], originally developed for sequence-to-sequence tasks in Natural Language Processing (NLP) [5]. The Transformer architecture has revolutionized NLP by enabling models to understand context through self-attention mechanisms. This capability helps the models to learn to dynamically assign weights to different parts of the input data to produce more accurate and contextually relevant outputs. The Transformer architecture has also been successfully applied to the field of chemistry in various tasks, such as single-step chemical reaction prediction [6, 7, 8], retrosynthesis [8, 9], molecule generation [10, 11, 12], and molecular property prediction [13, 14, 15]. However, these models are highly dependent on the quality and diversity of the data used for training.
Molecular data can be represented in the form of strings, which are derived from a graph traversal of their molecular structure. This compact representation, known as the Simplified Molecular Input Line Entry System (SMILES) [16] encodes molecules where each character represents an atom or bond, providing a linear representation useful for computational purposes and database storage. When the graph traversal follows the specific order established by the notation, the resulting SMILES is canonical, and a molecule has only one canonical SMILES but can have multiple non-canonical SMILES representations.
@@
Publicly available reaction datasets are obtained through data extraction from patents published by the United States Patent and Trademark Office (USPTO) [18]. These reactions are represented using SMILES representation along with their reactants and reagents. Previous work has created a sub-dataset called USPTO MIT [19] by filtering it to 470,000 examples and splitting it into train/test/validation sets. USPTO MIT datasets are widely used in the literature.
Since USPTO datasets are derived from patents, they provide highly specific reactions tailored exclusively to particular reactant-product combinations. Furthermore, the chemical space explored using these datasets is limited to the information contained within the patented reactions. Consequently, we assume that relying solely on patents introduces bias during training, leaving a substantial portion of the chemical space unaccounted for. While previous work [8] has demonstrated great accuracy in predicting reaction outcomes using this dataset, the lack of real-world applications stemming from such models highlights a key limitation: models trained solely on the patented chemical space suffer from generalization issues and are unsuitable for practical applications [20].
@@
Reaction template datasets are often argued to be difficult to maintain due to the endless and unmanageable number of reactant-product combinations [7], leading many to favor template-free methods as the future of the field. Maintaining templates for highly specific reactions, such as those in USPTO, is undeniably challenging. In contrast, we argue that using generic reactions, as in the proposed BRS, preserves the benefits of reaction templates while avoiding the associated maintenance issues.
To explore the potential of the proposed BRS, we introduce ProPreT5, a T5-based model tailored for the reaction product prediction task. ProPreT5 is trained in both template-based and template-free settings using the standard USPTO MIT dataset as well as a newly constructed dataset based on the generic BRS reaction templates. The aim of this work is not to improve performance on USPTO MIT, but to propose a more flexible and realistic alternative. For this reason, we adopt an economical training setup and use USPTO MIT primarily as a sanity check. In contrast, our main focus lies in training and evaluating the model on the more challenging dataset derived from BRS. To support learning in this setting, we also propose a novel augmentation strategy for SMARTS templates, designed to enhance generalization from a limited number of generic reactions.
The contributions presented in this article are threefold:
  • We address the limitations of publicly available reaction datasets by introducing a novel generic reaction set BRS, allowing for a broader exploration of the chemical space.
  • We release ProPreT5, an improved and highly flexible T5-based model capable of generating accurate reaction products. To the best of our knowledge, it is also the first large language model capable of handling and applying SMARTS templates.
  • We propose the first augmentation strategy for SMARTS templates, using specialization, generalization, combinations, and permutations of atom patterns to introduce chemically consistent diversity and improve model robustness.
TABLE I: Generic Reaction Patterns from the Broad Reaction Set (BRS)
@@
[…]
Note: The greater-than signs (>>>>) separates reactants from products, while a dot (.) distinguishes individual molecules. #n\#n represents any atom with the atomic number $n$, and specific letters indicate particular chemical environments; :n:n is used to map and track specific subgraphs within the reaction. For more information on the SMARTS notation, refer to [17].
@@ II Related Work
Most of the recent literature on template-based approaches in reaction prediction has focused on graph-based models. [22, 23]. However, to the best of our knowledge, the potential of Transformer models for template-based sequence prediction has yet to be explored.
Template-free, sequence-based models have achieved impressive results on the USPTO MIT benchmark, holding the current state-of-the-art in the single-step product prediction task [8]. However, these models face significant challenges in terms of applicability. The USPTO MIT dataset, derived from patented reactions, captures only a fraction of the diversity and complexity found in real-world organic synthesis, leaving large regions of the chemical space unexplored. As a result, while template-free models perform well in generating reaction outcomes, they struggle to generalize to the discovery of novel molecules or the prediction of reaction outcomes that fall outside the dataset [20].
Notable advancements in this domain include Molecular Transformer [6], which used a smaller version of the base Transformers architecture [4]. The model was trained on both forward and backward tasks. This was the first application of Transformers to the single-step reaction prediction task. Augmented Transformer [7] used an extensive data augmentation on input SMILES and demonstrated that doing so increased the model’s generalization. Chemformer [8], on the other hand, took things even further by proposing a much larger BART [24] model with special pretraining. The pretraining included masking and data augmentation, making the model more robust and increasing prediction accuracy. While these template-free Transformer models have demonstrated remarkable success in reaction prediction, the exploration of template-based sequence-to-sequence models for this task remains an underexplored area, leaving room for further development in this domain.
The lack of real-world applications of existing reaction prediction models reveals a critical weakness: their inability to generalize beyond narrow, benchmark-specific patterns. This lack highlights the need for a dataset with generic reactions, offering a more versatile training ground for reaction prediction models. This would allow models to explore the entire chemical space, ultimately advancing the field beyond the limitations of current benchmarks.
@@ III Datasets
Three datasets are used in this work: the USPTO MIT dataset [19]; BRS-Base, a dataset constructed using the proposed BRS generic reaction templates, one of the key contributions of this work; and BRS-Aug, a SMARTS-augmented version of the base training dataset, generated using the novel augmentation strategy proposed in this study, a second major contribution. The USPTO MIT dataset contains reactions, along with the corresponding reactants, reagents, and products extracted from patents, represented in SMILES notation. The USPTO MIT dataset includes highly specific reaction templates, valid only for the exact combination of reactants, reagents, and products. It is divided into two subsets: USPTO MIT Mixed, which does not distinguish between reactants and reagents and is considered a slightly more challenging problem, and USPTO MIT Separated, which separates reagents from reactants.
In this study, we worked with the USPTO MIT Mixed dataset with the standard train/validation/test split of approximately 409K, 30K, and 40K examples, respectively.
@@ III-A Proposed Reaction Set: Broad Reaction Set
In this study, we introduce 20 new generic reaction templates, as shown in Table I. Ten of these reactions are constructive, and the other ten are destructive, represented in SMARTS notation. These reactions are inspired by those used in an evolutionary algorithm named EvoMol [26], which demonstrated efficiency in exploring the chemical space [27]. These reactions serve as foundational building blocks for constructing reaction prediction datasets similar to USPTO. Since these reactions can be applied to a wide range of molecules, they offer significant flexibility for use with publicly available or commercial molecular datasets.
Destructive reactions, the reverse of constructive reactions, enable a return to an earlier stage in the chemical space, offering the possibility to explore alternative pathways. Constructive and destructive reactions are symmetrical.
@@
For simplicity, the transformations defined in these reactions are currently limited to atoms C, N, O, and F. However, this does not imply that the reactions cannot be applied to molecules containing other atoms. It simply means that only the reactivity of these four atoms is considered, and transformations will occur exclusively between them within the molecule. Additionally, reactions constructing cycles up to size 8 were defined. This limitation is partly due to challenges in defining reactions that can generate cycles of arbitrary size using SMARTS notation, but also because cycles larger than size 8 are rare. Moreover, chemists have proposed metrics that try to assess the synthesizability of a molecular graph, such as SAScore [28], where large cycles are penalized.
@@
III-B Dataset Construction
To construct the dataset used in this study, reactants were randomly sampled from two publicly available sources: EVO10 [29], which enumerates all possible molecules with up to 10 atoms of C, N, O, F, or S; and ChEMBL34 [31], a large-scale dataset of bioactive, drug-like molecules. ChEMBL34 was filtered to retain only molecules containing the same atom types as EVO10, ensuring chemical consistency and relevance across sources. The dataset was created using molecules sampled from these two sources. For each reaction in the template set, a molecule was randomly selected and evaluated for compatibility with the input pattern. If the structure matched, the reaction was executed using RDKit [25], an open-source cheminformatics library. If the reaction produced valid products, a series of filtering steps was applied to eliminate unrealistic outcomes commonly introduced by generic BRS templates. First, only canonical SMILES were retained. Then, we applied the filtering criteria from [29]: one filter removed molecules containing substructures not found in real-world compound datasets such as ChEMBL and ZINC [30]; the other excluded products containing Generic Cyclic Features (GCF), molecular ring scaffolds not observed in those datasets. Products that passed these filters were deemed realistic. To promote structural diversity, we retained multiple products for the same reactant–reaction combination, allowing the model to learn that a single input may correspond to multiple plausible outcomes. This resulted in the generation of the BRS-Base dataset consisting of 220K training, 10K validation, and 10K test examples. Much larger datasets can also be generated by applying the BRS templates to additional molecular sources.
III-C Data Augmentation Strategy for SMARTS Templates
In this study, a new data augmentation strategy for SMARTS templates is presented. To the best of our knowledge, this is the first attempt to augment this notation. Acquiring high-quality reaction templates is challenging, and in such cases, as in many other machine learning domains, data augmentation becomes a valuable tool. While previous work has successfully applied augmentation techniques to molecular SMILES representations, leading to improved performance in template-free models [32, 7, 8], no such strategy has been proposed for SMARTS. Inspired by these efforts, we introduce a novel augmentation framework tailored to SMARTS templates. This is particularly important in our setting, where the number of base templates is limited to just 20. Without augmentation, the model may struggle to grasp the syntax and semantics of the SMARTS language from such a small and rigid set, potentially resulting in a model that merely memorizes 20 isolated input–output transformations. Our augmentation strategy addresses this limitation by introducing chemically sound diversity into the templates, promoting generalization across reactions, and encouraging the model to learn the underlying semantics of the transformations rather than surface-level patterns.
The transformations applied to the SMARTS templates fall into four main categories:
  • Specialization: In SMARTS, atoms can be represented by their atomic numbers using the # symbol. For example, #6 matches any carbon atom. We specialize these general representations by replacing them with specific forms such as C (a carbon atom outside of a ring) or c (a carbon atom within a ring). This transformation restricts the template to apply only to cyclic or non-cyclic structures, introducing meaningful variation.
  • Generalization: Conversely, specific atom types like C or c can be generalized back to their atomic number form, such as #6, allowing the template to match a broader range of reactants. While the actual chemical context may vary, this transformation preserves the previous reaction mechanism and increases the model’s exposure to syntactic diversity.
  • Permutation: Atoms listed within SMARTS brackets, e.g., [#7, #8, #6; h:1], can be permuted without changing the chemical meaning. We apply permutations both within atom groups and between bracketed groups. For example, [#6, #7, #8; h:2].[O, N, F, C:1] can be swapped to [O, N, F, C:1].[#7, #8, #6; h:2]. These permutations increase syntactic diversity while preserving reactivity.
  • Combination: We generate multiple combinations of atom classes within SMARTS patterns. For instance, starting from [#6, #7, #8; h:1], we can create variants such as [#8; h:1], [#6, #7; h:1], and others. These combinations further expand the expressiveness of the templates.
There are important considerations when applying these transformations. Except for permutation, whenever a transformation is applied to the input (left-hand side) template, the same transformation must also be applied to the output (right-hand side) template. This is essential because, in chemical reactions, the atoms present in the input molecules persist in the output molecules, and maintaining consistency ensures chemical validity. In contrast, permutation only alters the order of atoms or groups within the SMARTS pattern without changing their identity or role, so it does not require a corresponding change in the output template. Each augmented template must be validated using RDKit [25] to ensure chemical correctness. Using these augmentation strategies, the size of the training dataset is doubled: the original entries are retained, and their augmented template versions are added. This results in a training set of 440K examples, referred to as BRS-Aug.